OpenStudy (anonymous):

the sum of 2 numbers is 15. if you double the smaller number and subtract the result from the larger number you get 48. what are the 2 numbers.

OpenStudy (lgbasallote):

let x be one number and y the other sum of 2 numbers is 15 so x + y = 15 double the smaller number so 2x then subtract result from larger number so y - 2x it says the result is 48 so y - 2x = 48 so y = 48 + 2x if i substitute this y in x + y = 15 i will get x + 48 + 2x = 15 3x + 48 = 15 3x = 15 - 48 3x = -33 x = -11 so that means y = 15 -(-11) = 15 + 11 = 26 got it?